# Vision Language Model-based Testing of Industrial Autonomous Mobile Robots

### v1

@@ 4.2 Experiment Setup
RVSG Feedback and Memory Settings. Based on a pilot study, we set the stopping condition for the feedback trigger to terminate after four configuration optimizations for the same scenario description, resulting in a total of five similar but distinct configurations per scenario. The stopping condition for the memory trigger was set to terminate after generating 10 sets of scenarios, each consisting of one scenario description and its corresponding optimal human configuration. Among these, the scenarios with the best and worst objective value were identified and repeated 30 times for further analysis.
@@ 4.3 Evaluation Metrics
  • Scenario Description Diversity (SDD): measures the diversity of natural language descriptions generated for scenarios, defined as one minus the average similarity computed by Sentence Transformers [22]. Higher values indicate greater diversity in scenario descriptions.
@@ 5.1 Comparison between RVSG and $\textit{RVSG}_{R}$— RQ1
Requirement
Route Objective Metric Performance Metric Diversity Metric
@@
Collision
Avoidance
(DTO $\downarrow$)
Route_1 3.020 / 0.666 0.2 / 3.7 2.002 / 2.222 8.482 / 8.491 0.249 / 0.299 0.142 / 0.253 4.550 / 4.950 60.415 / 41.385
@@
Stability
(Jerk $\uparrow$)
Route_1 0.980 / 1.371 0.2 / 0.3 2.002 / 2.020 8.482 / 8.472 0.249 / 0.248 0.142 / 0.200 4.550 / 8.228 60.415 / 61.838
@@
Efficiency
(TRG $\uparrow$)
Route_1 20.433 / 21.024 0.2 / 0.4 2.002 / 2.249 8.482 / 8.477 0.249 / 0.350 0.142 / 0.183 4.550 / 9.461 60.415 / 53.401
@@
For Diversity Metric, for SDD and HTD, RVSG outperformed $\textit{RVSG}_{R}$ across all requirements and routes, showing greater diversity in scenario descriptions and high-level task descriptions. However, for SSD, RVSG outperformed $\textit{RVSG}_{R}$ in 7 of 15 cases (0 for R1, 3 for R2, 4 for R3). This suggests that RVSG’s strength lies in linguistic and task-level diversity rather than in producing diverse human-robot interactions. A possible reason is that $\textit{RVSG}_{R}$, resembling random search, explores the simulation space more broadly.
Conclusion for RQ1: RVSG outperformed $\textit{RVSG}_{R}$ in generating scenarios that more effectively violate safety and performance requirements, degrading robot performance, and enhancing linguistic and task-level diversity. However, RVSG’s ability to simulate diverse scenarios is limited, likely due to its more guided search strategy.
@@ 5.2 Assessing Robot Behavioral Variability and Instability with RVSG-Generated Scenarios— RQ2
We picked the best and worst scenarios from all the scenarios generated by RVSG. We executed each selected scenario 30 times to study its characteristics for variability in the robot’s behaviors. For Objective Metric and Performance Metric, we performed the Mann-Whitney U test and the Vargha and Delaney Â12 effect size to compare differences between all the best ($\textit{RVSG}_{B}$) and worst ($\textit{RVSG}_{W}$) scenarios. We also calculated the standard deviation (std) for these metrics to understand the variability in the robot’s behavior across repeated executions. For Diversity Metric, we calculated SSD across the 30 simulations for each scenario. Table 2 summarizes the results for different requirements. For R1, results showed no statistically significant difference between $\textit{RVSG}_{B}$ and $\textit{RVSG}_{W}$ in all metrics except for DTO, indicating similar robot performance. Nevertheless, for DTO, $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$, meaning it better violated R1. Moreover, $\textit{RVSG}_{B}$ consistently achieved higher std values for almost all metrics. This indicates less stable robot behaviors when using $\textit{RVSG}_{B}$ compared to $\textit{RVSG}_{W}$, suggesting that $\textit{RVSG}_{B}$ is more suitable for uncovering unexpected behaviors during testing. For R2, $\textit{RVSG}_{B}$ achieved a comparable performance as $\textit{RVSG}_{W}$ for five out of seven metrics (i.e., Jerk, TRG, RPC, PL, TNM), and for the other two metrics (DTO, CHC), $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$. Regarding variability, $\textit{RVSG}_{B}$ consistently achieved higher std values across all metrics, indicating less stable robot behaviors during testing. Similarly, for R3, $\textit{RVSG}_{B}$ achieved comparable or significantly better performance than $\textit{RVSG}_{W}$ across the metrics, while also leading the robot to exhibit more unstable behaviors. This consistent pattern indicates that the best scenarios identified by RVSG are more effective at exposing unexpected or unstable behaviors of the robot, making them more suitable for testing. Finally, the diversity results (i.e., SSD) show that $\textit{RVSG}_{B}$ performed more unstable than $\textit{RVSG}_{W}$ for R1 and R3, while for R2, $\textit{RVSG}_{B}$ is more stable than $\textit{RVSG}_{W}$. This variation may lie in the requirements, and further experiments are needed to confirm this hypothesis.
@@ 5.3 Studying the Effect of Navigation Routes in RVSG-Generated Scenarios— RQ3
We further investigated the impact of different routes on the performance of the test scenarios generated by RVSG with same settings as RQ2 (i.e., picking best and worst scenario and the same statistics). We summarize the key findings here, and all detailed results are available in our public repository [20].
@@ 6 Threats to Validity
Concerning conclusion validity concerns, to draw a reliable conclusion, we employed appropriate statistical tests and a rigorous analysis based on established guides. Regarding internal validity concerns related to parameters, we used the default parameters for the VLM from OpenAI and kept the same parameter settings for all approaches. Concerning constructive validity threat, we used human-aware navigation metrics from HuNavSim [9]. These metrics are comparable and commonly applied in robot navigation [27, 28]. To mitigate threats due to hallucinations and errors in VLM outputs, we manually reviewed a subset of the generated responses. External Validity concerns the generalizability of RVSG and the empirical evaluation. We employed three requirements and five robot navigation routes. The selected requirements represent key safety and functionality of the robot and are provided by PAL Robotics. The navigation routes are selected in consultation with PAL Robotics to reflect real-world operations and human-robot interactions. To address the simulator fidelity threat, we used Gazebo, a high-fidelity simulator that enables realistic modeling of robot dynamics, sensor noise, and human-agent interactions. Moreover, this simulator is used by PAL Robotics in their practice. Besides, to enhance the realism of human behaviors, we used HuNavSim [9], which integrates the social force model and diverse human reactions to robot presence.
@@ 7 Lessons Learned and Industrial Perspectives
Benefits of VLMs in Test Scenario Generation. VLMs offer several key benefits for automated scenario generation for AMR navigation testing. First, they can process multimodal inputs (e.g., visual and textual data) enabling them to generate environmental context-dependent scenarios (including human behaviors) that are logical for a given environment. For example, in a warehouse setting, a VLM can generate realistic human behaviors, such as workers picking items from shelves. Second, VLMs possess commonsense reasoning capabilities [6], allowing them to infer plausible risky human behaviors that could violate requirements. Third, their natural language understanding enables interpretation of high-level F&S requirements, ensuring that generated scenarios are focused on violating those requirements. Fourth, VLMs support diverse scenario generation, which enhances test coverage. With feedback and memory mechanisms, RVSG iteratively refines scenario quality, allowing learning from historical failures, adjusting human behavior configurations, and avoiding redundancy.
Practical Implications. The practical adoption of RVSG can facilitate the development and testing of industrial AMRs. RVSG enables cost-effective and scalable testing by automating the generation of realistic, requirement-targeted scenarios without significant human effort. For example, a tester can specify requirements in natural language, which can then directly guide test generation, reducing both the need for technical expertise in robot testing and the overall implementation effort for coding tests. VLMs are demonstrated for one testing application in RVSG; however, they can also support other testing activities, such as automated test oracle generation and regression testing. By interpreting high-level requirements and analyzing multimodal outputs, VLMs can help determine expected outcomes, thereby reducing manual effort to devise test oracles. Furthermore, VLMs can support AMR regression testing by automatically identifying updates in software across evolution and using them to guide regression testing (e.g., test case selection). Lastly, VLMs can support other AMR development phases, including requirements engineering (e.g., requirements identification), modeling and analysis (e.g., tools for automated creation of robot simulation models), and software development (e.g., automating code generation).
@@ 8 Related Works
Testing of Robotic Systems. Various robotic testing approaches have been proposed, e.g., simulation-based testing [29, 30, 31, 32] and formal verification [33, 34]. For example, Hutchison et al. [35] proposed a framework to automate the test generation and execution for industrial autonomy systems, including robots, and robotics-oriented libraries. Brito et al. [36] proposed an integration testing approach for robotic systems focusing on the publish/subscribe communication function and generating critical test scenarios guided by coverage criteria. To ensure the correct ROS operation, Santos et al. [32] presented a property-based testing approach for ROS system configurations, focusing on orchestration errors that occur when integrating different ROS components at runtime. Compared to the above methods, RVSG focuses on system-level testing of an industrial robot to generate scenarios including human behaviors in simulation. It uses a VLM to search for semantically meaningful and realistic scenarios that challenge the robot’s behavior to violate its functional and safety requirements.
Foundation Models-based Test Generation. Foundation models (FMs), e.g., LLMs and VLMs, have shown benefits in many areas. Among them, FM-based test generation offers an effective way to generating tests [37, 38, 39]. With techniques, e.g., prompting or fine-tuning, FMs support testing activities, including unit testing [39, 40] and test oracle generation [41, 42]. Though FMs showed good performance in testing tasks like unit testing, their use to test complex software systems, e.g., robots, remains unexplored due to their complexity (e.g., real-time constraints and multimodal inputs), which pose unique testing challenges. Some works address these challenges for autonomous vehicles with FMs [43, 44]. In contrast, RVSG generates test scenarios for an industrial robot operating in dynamic environments. Specifically, RVSG focuses on scenarios involving human-robot interactions, which are critical to both functionality and safety. Using VLMs, RVSG can synthesize realistic scenarios that expose potential violations of robots’ safety and functional requirements.
@@ 9 Conclusions and Future Work
We present RVSG, a VLM-based test scenario generation approach for robot navigation. Given a functional or safety requirement, RVSG generates scenarios violating the requirement. RVSG employs a VLM to extract visual features from an image describing a robot’s environment. Next, the extracted features are then used by the VLM to generate human behavior configurations along with the scenario description by iteratively querying the VLM with specifically designed prompts. The scenarios generated in each iteration are then executed in simulation, and the execution results are used as feedback for the next iteration to refine the scenarios. We evaluated RVSG in a warehouse environment and selected three requirements and five robot navigation routes for the experiment. The results show that RVSG can generate effective requirement-violating test scenarios that increase the variability of robot behaviors, helping reveal behavioral uncertainties. Meanwhile, the navigation routes affect the performance and stability of RVSG-generated scenarios. Future works include studying the generalizability of RVSG by conducting experiments with other robots. Also, we plan to investigate multi-objective RVSG that can violate multiple requirements simultaneously. Our open-source repository is now publicly available online [20].
@@ Acknowledgments
This work is supported by the Co-tester project (No. 314544) and the Co-evolver project (No. 286898/F20), funded by the Research Council of Norway. The work is also supported by the RoboSAPIENS project funded by the European Commission’s Horizon Europe programme under grant agreement number 101133807. Aitor Arrieta is part of the Software and Systems Engineering research group of Mondragon Unibertsitatea (IT1519-22), supported by the Department of Education, Universities and Research of the Basque Country.

### v2

@@ 4.2 Experiment Setup
RVSG Feedback and Memory Settings. Based on a pilot study, we set the stopping condition for the feedback trigger to terminate after four configuration optimizations for the same scenario description, resulting in five similar but distinct configurations per scenario. The stopping condition for the memory trigger was set to terminate after generating 10 sets of scenarios, each consisting of one scenario description and its corresponding optimal human configuration. Among these, the scenarios with the best and worst objective value were identified and repeated 30 times for further analysis.
@@ 4.3 Evaluation Metrics
  • Scenario Description Diversity (SDD): measures the diversity of natural language scenario descriptions, defined as one minus the average similarity computed by Sentence Transformers [34]. Higher values indicate greater diversity in scenario descriptions.
@@ 5.1 Comparison between RVSG and $\textit{RVSG}_{R}$— RQ1
Requirement (Objective) Route Objective Metric Performance Metric Diversity Metric
@@
Collision Avoidance (DTO $\downarrow$) Route_1 3.020 / 0.666 0.2 / 3.7 2.002 / 2.222 8.482 / 8.491 0.249 / 0.299 0.142 / 0.253 4.550 / 4.950 60.415 / 41.385
@@
Stability (Jerk $\uparrow$) Route_1 0.980 / 1.371 0.2 / 0.3 2.002 / 2.020 8.482 / 8.472 0.249 / 0.248 0.142 / 0.200 4.550 / 8.228 60.415 / 61.838
@@
Efficiency (TRG $\uparrow$) Route_1 20.433 / 21.024 0.2 / 0.4 2.002 / 2.249 8.482 / 8.477 0.249 / 0.350 0.142 / 0.183 4.550 / 9.461 60.415 / 53.401
@@
For Diversity Metric, for SDD and HTD, RVSG outperformed $\textit{RVSG}_{R}$ across all requirements and routes, showing greater diversity in scenario descriptions and high-level task descriptions. However, for SSD, RVSG outperformed $\textit{RVSG}_{R}$ in 7 of 15 cases (0 for R1, 3 for R2, 4 for R3). This suggests that RVSG’s strength lies in linguistic and task-level diversity rather than in producing diverse human-robot interactions. A possible reason is that $\textit{RVSG}_{R}$, resembling random search, explores the search space more broadly.
Conclusion for RQ1: RVSG outperformed $\textit{RVSG}_{R}$ in generating scenarios that violate safety and performance requirements, degrading robot performance, and enhancing linguistic and task-level diversity. However, RVSG’s ability to simulate diverse scenarios is limited, likely due to its more guided search strategy.
@@ 5.2 Assessing Robot Behavioral Variability and Instability with RVSG-Generated Scenarios — RQ2
We picked the best and worst scenarios from all the scenarios generated by RVSG. We executed each selected scenario 30 times to study its characteristics for variability in the robot’s behaviors. For Objective Metric and Performance Metric, we performed the Mann-Whitney U test and the Vargha and Delaney Â12 effect size to compare differences between all the best ($\textit{RVSG}_{B}$) and worst ($\textit{RVSG}_{W}$) scenarios. We also calculated the standard deviation (std) for these metrics to understand the variability in the robot’s behavior across repeated executions. For Diversity Metric, we calculated SSD across the 30 simulations for each scenario. Table 2 summarizes the results for different requirements. For R1, results showed no statistically significant difference between $\textit{RVSG}_{B}$ and $\textit{RVSG}_{W}$ in all metrics except for DTO, indicating similar robot performance. Nevertheless, for DTO, $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$, meaning it better violated R1. Moreover, $\textit{RVSG}_{B}$ consistently achieved higher std values for almost all metrics. This indicates less stable robot behaviors when using $\textit{RVSG}_{B}$ compared to $\textit{RVSG}_{W}$, suggesting that $\textit{RVSG}_{B}$ is more suitable for uncovering unexpected behaviors during testing. For R2, $\textit{RVSG}_{B}$ achieved a comparable performance as $\textit{RVSG}_{W}$ for five out of seven metrics (i.e., Jerk, TRG, RPC, PL, TNM), and for the other two metrics (DTO, CHC), $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$. Regarding variability, $\textit{RVSG}_{B}$ consistently achieved higher std values across all metrics, indicating less stable robot behaviors during testing. Similarly, for R3, $\textit{RVSG}_{B}$ achieved comparable or significantly better performance than $\textit{RVSG}_{W}$ across the metrics, while also leading the robot to exhibit more unstable behaviors. This consistent pattern indicates the best scenarios identified by RVSG are more effective at exposing robot’s unexpected or unstable behaviors, making them more suitable for testing. Finally, the diversity results (i.e., SSD) show that $\textit{RVSG}_{B}$ performed more unstably than $\textit{RVSG}_{W}$ for R1 and R3, while for R2, $\textit{RVSG}_{B}$ is more stable than $\textit{RVSG}_{W}$. This variation may lie in the requirements, and further experiments are needed to confirm this hypothesis.
@@ 5.3 Studying the Effect of Navigation Routes in RVSG-Generated Scenarios — RQ3
We further investigated the impact of different routes on the performance of the test scenarios generated by RVSG with the same settings as RQ2 (i.e., picking the best and worst scenarios and the same statistics). We summarize the key findings here, and all detailed results are available in our public repository [45].
@@ 6 Threats to Validity
Concerning conclusion validity concerns, to draw a reliable conclusion, we employed appropriate statistical tests and a rigorous analysis based on established guides. Regarding internal validity concerns related to parameters, we used the default parameters for the VLM from OpenAI and kept the same parameter settings for all approaches. We also repeated each scenario 30 times to mitigate randomness in the simulator and ensure stable results. Concerning constructive validity threat, we used human-aware navigation metrics from HuNavSim [33]. These metrics are comparable and commonly applied in robot navigation [41, 7]. To mitigate threats due to hallucinations and errors in VLM outputs, we manually reviewed a subset of the generated responses.
An External Validity threat in our evaluation relates to the generalizability of the results. The first threat concerns obtaining the same level of benefits observed in our evaluation, with additional requirements, environments, and routes using TIAGO OMNI Base. To this end, we used three requirements, five robot navigation routes, and one environment, all of which were selected in conjunction with PAL to reflect real-world operations and human-robot interactions. Nonetheless, more experiments are warranted in the future. The second is the applicability of RVSG to other AMRs. To this end, experiments with additional AMRs are needed. However, note that we used the industrial AMR in a real industrial setting provided by PAL. Third, to address the simulator fidelity threat, we used Gazebo, a high-fidelity simulator that enables realistic modeling of robot dynamics, sensor noise, and human-agent interactions. Moreover, this simulator is used by PAL in their practice. Besides, to enhance the realism of human behaviors, we used HuNavSim [33], which integrates the social force model and behavior trees, and diverse human reactions to robot presence.
@@ 7 Lessons Learned and Industrial Perspectives
Lesson 1 – Requirements-driven guidance improves scenario relevance and diversity: Compared to unguided generation, our requirement-driven VLM-based test generation method consistently led to scenarios that not only violated the intended system requirements but also exposed diverse instabilities in our industrial AMR behavior in a realistic manner. This highlights the importance of aligning VLMs with domain-specific requirements rather than relying solely on open-ended prompting.
Lesson 2 – VLM-generated human behaviors increase simulation realism and fidelity: While Gazebo and HuNavSim [33] provide accurate physical and behavioral simulations, our technique makes the scenarios realistic and challenging in terms of human behaviors, thanks to the integration of VLM-based methods. This underscores that when testing AMRs, achieving fidelity is not only about physical fidelity but also about contextual realism of simulated entities.
Lesson 3 – Feedback and memory mechanisms enhance test diversity: Integrating simulation feedback and historical memory into RVSG improved linguistic and task-level diversity, producing better scenarios over time. However, this guided search strategy reduced the coverage of rare but potentially critical edge cases, which unguided approaches sometimes revealed. A hybrid strategy that combines guided and exploratory generation is envisioned as a future research avenue.
Lesson 4 – Navigation routes critically shape test outcomes: Our evaluation showed that scenario effectiveness varies depending on different navigation routes. While straight routes yielded fewer violations, routes with environmental complexities (e.g., routes close to shelves) increased requirement violations and instability. Our discussion with practitioners reinforced this finding, as they emphasized that realistic routes often include such complex and constrained layouts, making route diversity a critical factor when generating test cases.
Lesson 5 – RVSG supports the MAPLE-K loop enabling self-adaptation in robots: PAL Robotics is implementing the MAPLE-K loop inside its AMRs to allow them to deal with unsafe situations with self-adaptation. The work presented in this paper was carried out in collaboration with engineers from PAL, resulting in a tool that they will use to support various phases of MAPLE-K, including Plan, Legitimate, and Knowledge. PAL will utilize this tool for AMRs as well as other types of robots they are developing, and consider this tool an important addition to their software engineering tool suite for self-adaptive robotics.
Lesson 6 – Industrial adoption requires balancing automation with expert oversight: Discussions with our industrial partners confirmed that, while RVSG substantially reduces manual effort, lightweight expert validation is still recommended to ensure no hallucination, confirming, this way, scenario feasibility. While VLM-based testing can scale scenario generation, deployment at scale may still require a human-in-the-loop role to ensure both trustworthiness and acceptance in practice.
@@ 8 Related Works
Testing of Robotic Systems. Various robotic testing approaches have been proposed, e.g., simulation-based testing [2, 40, 16, 37] and formal verification [19, 23]. For example, Hutchison et al. [12] proposed a framework to automate the test generation and execution for industrial autonomy systems, including robots, and robotics-oriented libraries. Brito et al. [8] proposed an integration testing approach for robotic systems focusing on the publish/subscribe communication function and generating critical test scenarios guided by coverage criteria. To ensure the correct ROS operation, Santos et al. [37] presented a property-based testing approach for ROS system configurations, focusing on orchestration errors that occur when integrating different ROS components at runtime. To support the modeling and verification of robotic applications, Miyazawa et al. [23] proposed RoboChart, a domain-specific modeling language based on UML. RoboChart supports modeling and verification of robotics applications through model checking and theorem proving. Compared to the above methods, RVSG focuses on system-level testing of an industrial robot to generate scenarios including human behaviors in simulation. It uses a VLM to search for semantically meaningful and realistic scenarios that challenge the robot’s behavior to violate its functional and safety requirements.
Foundation Models-based Test Generation. Foundation models (FMs), e.g., LLMs and VLMs, have shown benefits in many areas. Among them, FM-based test generation offers an effective way to generate tests [44, 11, 38]. With techniques, e.g., prompting or fine-tuning, FMs support testing activities, including unit testing [38, 42] and test oracle generation [43, 25]. For instance, Schäfer et al. [38] proposed TESTPILOT, an adaptive LLM-based test generation tool for JavaScript. TESTPILOT generates unit tests by iteratively querying an LLM with a specifically designed prompt. FMs also show promise in test minimization. For example, Pan et al. [31] proposed LTM, a test suite minimization approach based on LLMs and similarity measures. Though FMs showed good performance in testing tasks like unit testing, their use to test complex software systems, e.g., robots, remains unexplored due to their complexity (e.g., real-time constraints and multimodal inputs), which pose unique testing challenges. Some works address these challenges for autonomous vehicles with FMs [18, 17]. In contrast, RVSG generates test scenarios for an industrial robot operating in dynamic environments. Specifically, RVSG focuses on scenarios involving human-robot interactions, which are critical to both functionality and safety. Using VLMs, RVSG can synthesize realistic scenarios that expose potential violations of robots’ safety and functional requirements.
@@ 9 Conclusions and Future Work
We present RVSG, a VLM-based test scenario generation approach for robot navigation. Given a functional or safety requirement, RVSG generates scenarios violating the requirement. RVSG employs a VLM to extract visual features from an image describing a robot’s environment. Next, the extracted features are then used by the VLM to generate human behavior configurations along with the scenario description by iteratively querying the VLM with specifically designed prompts. The scenarios generated in each iteration are then executed in simulation, and the execution results are used as feedback for the next iteration to refine the scenarios. We evaluated RVSG in a warehouse environment and selected three requirements and five robot navigation routes for the experiment. The results show that RVSG can generate effective requirement-violating test scenarios that increase the variability of robot behaviors, helping reveal behavioral uncertainties. Meanwhile, the navigation routes affect the performance and stability of RVSG-generated scenarios. Future works include studying the generalizability of RVSG by conducting experiments with other robots. Also, we plan to investigate multi-objective RVSG that can violate multiple requirements simultaneously. Our repository is now publicly available online [45].
@@ Acknowledgments
This work is supported by the RoboSAPIENS project funded by the European Commission’s Horizon Europe programme (grant No. 101133807), and by the Co-tester (No. 314544) and Co-evolver (No. 286898/F20) projects funded by the Research Council of Norway. Aitor Arrieta is part of the Software and Systems Engineering research group of Mondragon Unibertsitatea (IT1519-22), supported by the Department of Education, Universities and Research of the Basque Country. Aitor Arrieta is also supported by the Spanish Ministry of Science, Innovation and Universities (project PID2023-152979OA-I00), funded by MCIU /AEI /10.13039/501100011033 / FEDER, UE.